Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by f2001290 » Tue May 29, 2007 7:30 am
A Four digit safe code does not contain the digits 1 and 4 at all. What is the probability that it has at least one even digit?

I don't have an OA.

P(atleast one digit is even) = 1 - P(all digit odds)

We have 4 odd digits excluding 1 (3,5,7,9).
So, possible 4 digit odd numbers = 4*4*4*4
All possible 4 digit numbers = 7*8*8*8 (Ensure that 1st digit is not 0 and exclude 1 and 4)

P(all digits odd) = 4*4*4*4/7*8*8*8 = 1/14
P(atleast one digit is even) = 1 - 1/14 = 13/14

Any different solution for this . I am very bad at probability
Join the discussion
Source: — Problem Solving |

by RAGS » Tue May 29, 2007 11:03 am
4 digit safe code without 1 and 4
possible digits for each digit 0,2,3,5,6,7,8,9 ie 8 ways
total 8*8*8*8

consider all odd digits so possible digits 3,5,7,9 ie 4 ways
total 4*4*4*4

So atleast one even is= 8*8*8*8 -4*4*4*4 / 8*8*8*8

= 1 - 1/16 = 15/16

Why should the first digit not be zero????????????
If It Is To Be It Is Up To Me
Join the discussion

by f2001290 » Tue May 29, 2007 10:18 pm
If 1st digit is zero, then it becomes a 3 digit number.

0123 is a 3 digit number.

Rags, I may be wrong (most of the time) :wink: . The most unfortunate thing is that I don't have an OA
Join the discussion

by gabriel » Wed May 30, 2007 8:46 am
... i agree with RAGS ... 0 can be the first number .. wen u say that 0 as the first number wuld reduce it to a 3 digit number .. that is only true wen u consider the numerical value of the 4 digit number ... over here the numerical value is not important but the combination of different numbers is.. so 0 can be the first number ... btw my bag which had a number lock used to open with a 0000 :D
Join the discussion